OpenStudy (rissyroo118):

Medal will be given :D 4x - 3y = k -8x + 6y = 10 If the pair of equations has infinitely many solutions, what is the value of k? A)-10 B)-5 C)5 D)10

OpenStudy (anonymous):

Lets start by just leaving k in there and trying to combine the two equations to solve. Can you see an easy way to combine the 2 equations into 1?

OpenStudy (rissyroo118):

i multiple , but i'd say k=10/-2=-5 so the answer i come up with is B

OpenStudy (anonymous):

right so (i think) the easiest way is to add 2 times the first equation to the second equation: 2*4x-2*3y+(-8x+6y)=2*k+10 left hand side goes to zero, 2k+10=0 k=-5 you're absolutely right!